OpenStudy (anonymous):

determine the lowest unit price. brand a: 29 ox for $8.99 Brand b: 2lb, 8 oz for $11.60

OpenStudy (anonymous):

8.99/29 = .31 11.60/40 = .29 Brand b: 2lb, 8 oz for $11.60

OpenStudy (anonymous):

how many ounces in a pound?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

which makes brand B 40 ounces

OpenStudy (anonymous):

So set up a proportion.

OpenStudy (anonymous):

\[\frac{8.99}{29}=unit price \frac{11.60}{40}=unit price\]

OpenStudy (anonymous):

So it's .31 (aka 31 cents) per ounce for brand a, and .29 (29 cents) per pound for brand b

OpenStudy (anonymous):

so brand B is the better buy

OpenStudy (anonymous):

oh okay, great answer your brain was working good. thanks a lot

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

joan can mow a 10-acre field in 5 hours. How long would it take her to moa a 4-acre field?